# Liouville phenomenon for the Klein-Gordon equation in $1+1$ dimensions

### 04 Mar 2026 v1

@@ 7.2. Checking the solution formula
(7.4) $\partial_{x}\partial_{y}u(x,y)=\partial_{x}\partial_{y}(f(x)+g(y))-f(0)\partial_{x}\partial_{y}J_{0,0}(x,y)\\ -\partial_{x}\partial_{y}\int_{0}^{x}J_{1,0}(y,x-s)f(s)\,\mathrm{d}s-\partial_{x}\partial_{y}\int_{0}^{y}J_{1,0}(x,y-t)g(t)\,\mathrm{d}t\\ =f(0)J_{0,0}(x,y)-\partial_{x}\partial_{y}\int_{0}^{x}J_{1,0}(y,x-s)f(s)\,\mathrm{d}s-\partial_{x}\partial_{y}\int_{0}^{y}J_{1,0}(x,y-t)g(t)\,\mathrm{d}t,$
@@
(7.5) $\partial_{x}\int_{0}^{x}J_{1,0}(y,x-s)f(s)\,\mathrm{d}s=J_{1,0}(y,0)f(x)+\int_{0}^{x}\partial_{x}J_{1,0}(y,x-s)f(s)\,\mathrm{d}s\\ =f(x)y+\int_{0}^{x}\partial_{x}J_{1,0}(y,x-s)f(s)\,\mathrm{d}s,$
@@
(7.6) $\partial_{x}\partial_{y}\int_{0}^{x}J_{1,0}(y,x-s)f(s)\,\mathrm{d}s=\\ =f(x)+\int_{0}^{x}\partial_{x}\partial_{y}J_{1,0}(y,x-s)f(s)\,\mathrm{d}s=f(x)-\int_{0}^{x}J_{1,0}(y,x-s)f(s)\,\mathrm{d}s.$
@@ Acknowledgements
The second author acknowledges support from Vetenskapsrådet (VR grant 2020-03733).

### 25 Mar 2026 v3

@@ 7.2. Checking the solution formula
(7.4) $\partial_{x}\partial_{y}u(x,y)=\partial_{x}\partial_{y}(f(x)+g(y))-f(0)\partial_{x}\partial_{y}J_{0,0}(x,y)\\ -\partial_{x}\partial_{y}\int_{0}^{x}J_{0,1}(x-s,y)f(s)\,\mathrm{d}s-\partial_{x}\partial_{y}\int_{0}^{y}J_{1,0}(x,y-t)g(t)\,\mathrm{d}t\\ =f(0)J_{0,0}(x,y)-\partial_{x}\partial_{y}\int_{0}^{x}J_{0,1}(x-s,y)f(s)\,\mathrm{d}s-\partial_{x}\partial_{y}\int_{0}^{y}J_{1,0}(x,y-t)g(t)\,\mathrm{d}t,$
@@
(7.5) $\partial_{x}\int_{0}^{x}J_{0,1}(x-s,y)f(s)\,\mathrm{d}s=J_{0,1}(0,y)f(x)+\int_{0}^{x}\partial_{x}J_{0,1}(x-s,y)f(s)\,\mathrm{d}s\\ =f(x)y+\int_{0}^{x}\partial_{x}J_{0,1}(x-s,y)f(s)\,\mathrm{d}s,$
@@
(7.6) $\partial_{x}\partial_{y}\int_{0}^{x}J_{0,1}(x-s,y)f(s)\,\mathrm{d}s=\\ =f(x)+\int_{0}^{x}\partial_{x}\partial_{y}J_{0,1}(x-s,y)f(s)\,\mathrm{d}s=f(x)-\int_{0}^{x}J_{0,1}(x-s,y)f(s)\,\mathrm{d}s.$
@@ Acknowledgements
The author acknowledges support from Vetenskapsrådet (VR grant 2020-03733), and by grant 075-15-2021-602 of the Government of the Russian Federation for the state support of scientific research, carried out under the supervision of leading scientists.